OpenStudy (anonymous):

Summarize all pertinent information and sketch the graph of y = f(x): f(x) = sqrt (25-x^2)

OpenStudy (anonymous):

Anyone there?

OpenStudy (anonymous):

Well, so what is pertinent? Your square root argument must not be negative, which restricts your domain

OpenStudy (anonymous):

so -5<=x<=5

OpenStudy (anonymous):

Next, how does that restrict the domain? x^2 is always positive, so the biggest the square root part can be is 5, and the smallest is 0, so the range is 0<=y<=5

OpenStudy (anonymous):

I don't know what your teacher means by "pertinent information," but you can graph from there.

OpenStudy (anonymous):

Square both sides of the equation y=(25-x^2) and rearrange the terms as shown below: \[y^2 + x^2=5^2\] I might be wrong but isn't that a circle, centered at the origin with a radius of 5?

OpenStudy (anonymous):

but the circle is not a function. You often can't just square both sides when it is a function. If the f(x) was really a y, then yes, it is a circle radius 5

OpenStudy (anonymous):

What we actually get looks to be the top half of that circle